HUBBLE SPACE TELESCOPE DAILY REPORT #1667 PERIOD COVERED: 0000Z (UTC) 07/01/96 - 0000Z (UTC) 07/02/96 Daily Status Report as of 184/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6364 (Snapshot Survey of Proto-Planetary Nebulae = and AGB Stars) The WF/PC-2 was used to make some snapshot survey observations of = protoplanetary nebular system R-Crb to study its circumstellar envelope. = The purpose is to discern the evolutionary sequence that produces = axially-symmetric planetary nebulae. The observations were executed as = scheduled, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 6188 (WF/PC-2 Cycle 5 Calibration: = Dark) The WF/PC-2 was used to make internal dark calibration images. = The observations were completed as scheduled, and no problems were = reported. 1.3 Completed FOC/96 5835 (The Structure of the Inner Coma of the = Comet Chiron) The Faint Object Camera (f/96) was used to make observations of = the object/comet 2060 Chiron in order to study its inner coma. The goals = are to map the structure and color of the inner coma. The observations = were executed as planned, and no problems were reported. 1.4 Completed WF/PC-2 6678 (Nearby Young Dwarf Galaxies and Their = Lyman Alpha Emissions) The WF/PC-2 was used to make observations of the young, metal = deficient, blue, compact, dwarf galaxy T1214-277 in order to study its = Lyman Alpha emissions. The observations were completed as planned, and = no problems were reported. 1.5 Completed WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were = reported. 1.6 Completed FOS/BL 6612 (Damped Lyman-Alpha Absorption Lines From = Moderate Redshifted Galaxies) The Faint Object Spectrograph (Blue Side) was used to make = observations of the moderate redshifted galaxy/QSO UM301 in order to = study the damped Lyman-Alpha absorption lines which are located in the = line of sight to the object. In addition, the neutral H column density = will be determined. The observations were completed as planed and no = problems were reported. 1.7 Completed Two Sets of WF/PC-2 6187 (WF/PC-2 Cycle 5 Calibration: = Earth Flat) The WF/PC-2 was used to make Earth streak flat calibration images. = These streak flats will be used to produce flat fields, to map the OTA = illumination pattern and, in combination with other data, to produce new = superflats. The observations were completed as planned, and no problems = were reported. 1.8 Completed FOS/RD 6059 (Imaging Low-Mass Companions of Nearby = Stars) The Faint Object Spectrograph (Red Side) was used to make = observations of low-mass nearby star GL293 to search for possible faint = companions. The bright star was placed behind an occulting bar. The = observations were executed as scheduled, and no anomalies were noted. 1.9 Completed GHRS/1/2 6594 (White Dwarf Companions of Barium and CH = Peculiar Stars) The High Resolution Spectrograph (Sides One and Two) were used to = attempt to make spectrographic observations of the white dwarf of the = Barium and CH Peculiar class star HD104979. This will test the theory = that abundance anomalies in this star caused by mass transfer from the pre= cursor of the white dwarf. The observations were executed as scheduled, = and no anomalies were noted. 1.10 Completed FOS/RD 6167 (FOS Cycle 5: Dark Monitoring) The Faint Object Spectrograph (Red Side) was used to make several = dark calibration images. All exposures were internals. All spectra was = done in the image mode. The observations were completed as planned, and = no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 16 Successful : 15 The update at 184/0354Z failed. 2.3 Operations Notes: HST operations transferred to the SAMS backup facility at = 184/0700Z in preparation for today's simulation. 3.0 SIGNIFICANT FORTHCOMING EVENTS: GSFC Internal Simulation #7 (Deploy) will be conducted today from = 8:00 a.m. until 4:00 p.m. /DMH